. A company issued 5-year, 7% bonds with a par value of $250,000. The market rate when the bonds were issued was 4.0% on 1/1/19. Interest is paid 2 times per year on 6.30 and 12.31. Calculate the sale price and record the journal entry for this sale. Using the straight-line method, calculate the amount of interest expense for the first semiannual interest period and record the journal entry
Answer #1
Determination of Sale Price :-
Face Value of Bond = $250,000
Life of Bond (in years) = 5
Maturity amount of Bond $250,000
Coupon rate = 7% per year payable semi annually i.e. 3.5% per semi annual
Semi annual Coupon Amount = $250,000 × 3.5%
= $8,750 This will be received 5*2 =10 times
n = 5 years i.e. 5 * 2 = 10 semi annuals
Market rate on similar bonds (r) = 4% p.a.
= 2% semi annual
PVAF(r, n) = Present Value Annuity factor at r% for n periods
Present Value of an Annuity of $1 at 2% for 10 periods = 8.98259
PVIF(r, n) = Present Value interest factor at r% for n periods
Present Value of 1 at 2% at n= 10 = 0.82035
Sale Price of Bond = Coupon Amount * PVAF(r,n)     + Redemption Amount * PVIF(r,n)
= $8750 * 8.98259 + $250000 * 0.82035
= $78,598 + $205,088
= $283,685
Date Account Titles and Explanation Debit Credit
01/01/2019 Cash $                                   283,685
              Bonds Payable $             250,000
              Premium on bonds payable $               33,685
(To record sale of bonds on premium)
Determination of amount of interest expense for the first semiannual interest period :-
Straight Line method of Bond Amortization:-
Annual straight-line amortization of the bond premium = 33685 / 5 years
= $6,737
Semi-Annual straight-line amortization of the bond premium = $6737/2
= $3,369
Semi annual Coupon Amount = $8,750
Less: Semi annual amortization of the bond premium = $3,369
Interest expense for the first semiannual interest periodn = $5,381
Date Account Titles and Explanation Debit Credit
30/06/2019 Interest Expense $                                       5,381
Premium on bonds payable $                                       3,369
              Cash ($250,000 * 7% * 6/12) $                 8,750
(To record the payment of interest)
31/12/2019 Interest Expense $                                       5,381
Premium on bonds payable $                                       3,369
              Cash ($250,000 * 7% * 6/12) $                 8,750
(To record the payment of interest)

Feel free to ask any clarification, if required. Please provide feedback by thumbs up, if satisfied. It will be highly appreciated. Thank you.